sin((theta divided by 3) + (pie divided by 4)) = one half
1
Expert's answer
2012-01-06T08:27:17-0500
sin((theta divided by 3) + (pie divided by 4)) = one half
theta/3
+pie/4=pie/6 +2*pie*k or theta/3 +pie/4=5*pie/6 +2*pie*k, where k is
integer
then theta/3=pie/6 -pie/4 +2*pie*k or theta/3=5*pie/6 -pie/4
+2*pie*k, where k is integer
then theta/3= -pie/12 +2*pie*k or
theta/3=7*pie/12 +2*pie*k, where k is integer
then theta= -pie/4 +6*pie*k or
theta=7*pie/4 +6*pie*k, where k is integer